OpenStudy (anonymous):

Solve |2x + 8| = 20.

OpenStudy (karatechopper):

ok in this problem there are also two cases. 2x+8=20 and -20=2x+8

OpenStudy (karatechopper):

Now it must be easy to solve!

OpenStudy (anonymous):

x=6

OpenStudy (karatechopper):

we can start with 2x+8=20 first 2x+8=20 2x+8-8=20-8 2x=12 2x/2=12/2 x=6 lets to the second now! -20=2x+8 -20-8=2x+8-8 -28=2x -28/2=2x/2 -14=x

OpenStudy (anonymous):

ur such good help thnx

OpenStudy (karatechopper):

most welcome:) just ask if need help in more